# InvFussion: Bridging Supervised and Zero-shot Diffusion for Inverse Problems

### v1

@@ A.1 Architecture and Model Training
We plan to publish our code implementation and model checkpoints upon acceptance.
@@
Models are trained and evaluated with the official implementation of the EDM2 [29] training script, using the default P_mean = -0.8, P_std = 1.6, float16 mixed precision, and a learning rate of $55e-$ 5, using the default learning rate scheduler. Sampling is done with the default Heun scheduler using a total of NFEs for sampling. Additional experiment specific-training hyperparameters can be found in Tab. 6. The augmentation implementation and hyperparameters are cloned from EDM [28].
A.2 Degradations
A.2.1 Patch In-Painting
Patch in-painting is implemented as any in-painting mask where $p\in(0,0.1)$ of patches of some unvarying size are visible. When creating a new mask, the patch size, $p$, and visible patches are randomly sampled until a valid mask in which some patches are visible is created. We preform this “rejection sampling” to remove masks in which no patches are visible.
A.2.2 Strided Motion-Blur
This family of degradations implements a strided convolution operator, for which we specifically choose a motion-blur kernel. To create a new degradation operator, we first sample a motion-blur kernel the size of the input image using the motionblur library [^3] https://github.com/LeviBorodenko/motionblur. We then sample a stride size from $[2,4,8]$ and smooth the motion-blur kernel by convolving it with a isometric Gaussian kernel with standard deviation equal to the stride. The kernels are normalized for numeric stability.
A.2.3 Matrix Degradation
The matrix degradation is a matrix operator with $(1,128)$ rows, each of which is sampled from a multivariate normal distribution and normalized. The number of rows is limited to enable the use of different matrices across the batch dimension, which is more computationally intensive. The pseudo-inverse operator is computed directly with pytorch’s p_inv function.
A.2.4 Box Out-Painting
A.3 Zero-Shot Methods
We use our own implementations for all zero-shot methods, using the default hyperparameter unless specified otherwise. We use the variance-exploding notation, and we adjust all sampling algorithms accordingly. DPS [9] and DAPS [56] only require access to the degradation operator ${\bm{H}}$, while DDNM [55] and $\Pi$GDM [47] also make use of the pseudo-inverse operator ${\bm{H}}^{\dagger}$. In the original implementation DDRM [31] makes use of the complete SVD of the operator ${\bm{H}}$. Because obtaining this SVD is computationally expensive (or even practically impossible in reasonable time) for most of our degradations, we instead implement the DDRM algorithm using only ${\bm{H}}$ and ${\bm{H}}^{\dagger}$. For the noise-less case, this implementation is equivalent to the original implementation. This is not true case in the general noisy case, unless all the degradation’s singular values that are not zeros are equal. Due to the degradation families we choose to apply, we assume that most of the singular values that are not zeros are nearly one for the purpose of implementing DDRM in the noisy degradation experiment. For $\Pi$GDM, we make a similar assumption to implement the noisy sampling algorithm, to avoid inverting a large matrix. For the in-painting case, the assumption holds. Similarly, in the strided motion-blur we make use of the fact that white Gaussian noise remains white Gaussian noise under a Fourier transform, and that the application of the kernel is multiplicative.
@@ Appendix B Using Diffusion Models for MMSE
In Sec. 3.1 we explore the use of the InvFussion architecture of MMSE estimation. A DNN can be trained directly as a degradation-aware MMSE estimator using a regression loss as seen in Eq. 9. Nevertheless, we notice that we can utilize our existing trained diffusion models (trained using Eq. 8) for the same task. Intuitively, this is because for sufficiently high noise values, the noisy input to the network is equivalent to pure noise. Therefore, The network learns to rely only on the information from ${\mathbf{y}}$ and ${\bm{H}}$. In practice, we use the value $\sigma_{t}=100$ for the MMSE estimation.
We test the gap between models trained directly for MMSE estimation (Eq. 9) and models trained for diffusion (Eq. 8) on FFHQ64, using the same architecture. The MMSE model’s input is Gaussian noise, instead of the noisy image input to the diffusion model, to keep the architecture identical. The results in Tab. 7 reflect that both methods yield approximately the same results, whether trained using the degradation-aware InvFussion architecture and loss or without it. We conclude that it is probably more cost-effective to use a single model for both posterior sampling and MMSE estimation task, and do so for the experiments in our paper.
@@
Trained for MMSE (Eq. 9)
Palette [42] $23.08$ $19.36$ $15.63$
InvFussion $24.66$ $19.79$ $17.92$
Trained For diffusion (Eq. 8)
Palette [42] $23.47$ $19.65$ $15.51$
InvFussion $24.74$ $19.81$ $17.62$
Appendix C Additional Results
C.1 Noisy Degradations
Table 8 shows the results of the experiment described in Sec. 4.2. The results show that InvFussion maintains superiority when noise is added to the degradation. We expect this framework to work well across many noise types, not limited to white Gaussian noise, as the model learns to treat the noise through training, which does not rely on any fundumental properties of the noise.
@@ C.1 Noisy Degradations
DDRM [31] 11.69 137.04 12.21 155.59 11.80 86.72 $25$
DDNM [55] 19.82 97.22 17.26 46.44 11.94 96.52 $100$
DPS [9] 6.98 12.34 8.26 7.77 9.18 27.33 $1000$*
$\Pi$GDM [47] 16.49 11.50 15.69 21.56 10.76 10.91 $100$*
DAPS [56] 16.90 46.41 15.54 43.72 10.57 58.67 $1000$
@@
Palette [42] 17.09 6.05 16.70 5.21 12.58 6.59 $63$
InvFussion 18.09 5.39 16.81 4.57 14.70 6.51 $63$
MMSE estimation methods (Training-based)
Palette [42] 19.82 - 18.89 - 15.59 - $1$
InvFussion 20.67 - 19.61 - 17.48 - $1$
C.2 FFHQ 256
Table 9 shows quantitative results for our experiment on $256\times 256$ FFHQ images. InvFussion remains the SOTA method as measured by C-FID – the perceptual image quality of valid solutions to the inverse problem.
@@ C.2 FFHQ 256
DDRM [31] $25.33$ $45.21$ $45.22$ $12.76$ $103.96$ $103.96$ $25$
DDNM [55] $26.21$ $51.19$ $51.15$ $14.25$ $105.51$ $105.51$ $100$
DPS [9] $22.70$ $18.32$ $22.42$ $9.71$ $26.67$ $94.34$ $1000$*
$\Pi$GDM [47] $24.63$ $11.81$ $11.82$ $16.66$ $26.63$ $26.63$ $100$*
DAPS [56] $15.11$ $286.76$ $96.40$ $16.99$ $44.96$ $50.59$ $1000$
@@
Palette [42] $23.20$ $8.67$ $11.86$ $18.22$ $10.65$ $11.89$ $63$
InvFussion $23.75$ $9.52$ $10.70$ $18.54$ $10.64$ $10.78$ $63$
C.3 Unconditional Samples
Figures 6 and 7 show qualitative examples for unconditional samples from our model i.e. samples that are not conditioned on a measurement vector ${\mathbf{y}}$. The models trained on ImageNet are conditioned on the class label, and use CFG [25] coefficient of 2.0 to enhance the image quality.
Figure 6: Examples of unconditional samples from FFHQ.
Figure 7: Examples of class-conditional samples from ImageNet. Samples are not conditioned on any measurements ${\mathbf{y}}$, and are sampled with CFG of 2.0 on the class conditioning.

### v2

@@ A.1 Architecture and Model Training
Code implementation available at https://github.com/noamelata/InvFusion.
Our experiments use data from the FFHQ [ffhq] (CC-BY 4.0 license) and ImageNet [imagenet] datasets. We implement all models using the official implementation of HDiT [crowson2024scalable] ImageTransformerDenoiserModelV2 architecture (MIT license). We use default hyperparameters, with changes listed in Tab.˜5. The initial patch size is indicated in the table, with the patch size doubling with the progression along the list shown in the ‘Depths’ column. The attention type column signified which type of attention was used. ‘NAttnxx’ indicates neighborhood attention [hassani2023neighborhood, hassani2024faster] with a kernel of size $x$, which performs attention between patches only in an $x$ sized neighborhood. ‘GAttn’ indicates global attention. As described in Sec.˜3, InvFusion requires that all layer widths are divisible by the input channels, which in this case is $3$. Also each layer’s width times the layer’s feature map resolution must be a multiple of the input dimension $D$. In our experiments, this held for all but the deepest layers in the FFHQ256 model. For this reason, we did not apply InvFusion to these layers, and used the baseline unconditional attention implementation instead. This provides evidence that our InvFusion architecture works even when not applied evenly on all the network’s attention layers. In preliminary experiments, we found joint-attention, in which keys and values from the feature degradation layer are concatenated with the keys and value in the main (self) stream, is preferable to applying self-attention followed by cross-attention.
Models are trained and evaluated with the official implementation of the EDM2 [karras2023analyzing] training script (licensed as CC BY-NC-SA 4.0), using the default P_mean = -0.8, P_std = 1.6, and a learning rate of 5e55e-5, using the default learning rate scheduler. Sampling is done with the default Heun scheduler using a total of $63$ NFEs for sampling. Additional experiment specific-training hyperparameters can be found in Tab.˜6. The augmentation implementation and hyperparameters are cloned from EDM [karras2022elucidating]. We have used 8 Nvidia A40 GPUs (or equivalent hardware) with 49GB of memory for all experiments. Training our model takes approximately one day for the FFHQ64 models, two days for the FFHQ256 model, and four days for the ImageNet64 model.
A.2 Pseudo-Code
A.2.1 Feature Degradation Layer
The Feature Degradation Layer is implemented with the following steps, as described below; The activations are un-patched into the image shape, after which the degradation is applied to them. Next, the measurements are concatenated to the degraded features, a linear function and an activation are applied, and the pseudo-inverse degradation takes the degraded features back into the images space.
1class FeatureDegradation($n$.Module):
2 def __init__(self, channels, patch_size, im_channels=3):
3 super().__init__()
4 self.h = patch_size[0]
5 self.w = patch_size[1]
6 self.im_channels = im_channels
7 self.deg_linear = Linear(channels + 1, channels, bias=True)
8
9 def forward(self, x, degradation, y):
10 x = rearrange(x,
11 "... h w (nh nw k c) -> ... k c (h nh) (w nw)",
12 nh=self.h, nw=self.w, c=self.im_channels)
13 _y = degradation.H(x)
14 _y = torch.cat([y, _y], -2)
15 _y = act(self.deg_linear(_y))
16 _x = degradation.H_pinv(_y)
17 _x = rearrange(_x,
18 "... k c (h nh) (w nw) -> ... h w (nh nw k c)",
19 nh=self.h, nw=self.w)
20 return _x
A.2.2 InvFusion Block
Below is the implementation of an InvFusion block applied on global attention. The input $\mathbf{x}$} is projected to create queries, values and keys, written as ${\mathbf{q}},{\mathbf{k}},{\mathbf{v}}$ respectively. The conditional input $\mathbf{y}$}, which is the output of the Feature Degradation layer, is also projected to make additional keys and values, ${\mathbf{k}}^{\prime},{\mathbf{v}}^{\prime}$, which are concatenated to the previous keys and values before the attention is computed. We find that projecting both $\mathbf{x}$} and $\mathbf{y}$} to produce $\mathbf{k}$$^{\prime}$ is beneficial, as the query key matching may be harder to learn solely on the outputs of the Feature Degradation layers. The neighborhood attention of an InvFusion block is similar, using the native neighborhood attention [hassani2023neighborhood, hassani2024faster] implementation to compute the attention itself.
1class SelfAttentionBlock($n$.Module):
2 def __init__(self, d, d_head, cond, dropout=0.0,
3 joint=False # Whether to apply InvFusion in this layer
4 ):
5 super().__init__()
6 self.d_head = d_head
7 self.n_heads = d // d_head
8 self.norm = AdaRMSNorm(d, cond)
9 self.jnorm = AdaRMSNorm(d, cond) if joint else None
10 self.qkv_proj = Linear(d, d * 3, bias=False)
11 self.jk_proj = Linear(d * 2, d, bias=False) if joint else None
12 self.jv_proj = Linear(d, d, bias=False) if joint else None
13 self.scale = $n$.Parameter(torch.full([self.n_heads], 10.0))
14 self.pos_emb = AxialRoPE(d_head // 2, self.n_heads)
15 self.dropout = $n$.Dropout(dropout)
16 self.out_proj = zero_init(Linear(d, d, bias=False))
17
18 def forward(self, x, theta, cond, y=None):
19 skip = x
20 x = self.norm(x, cond)
21 qkv = self.qkv_proj(x)
22 q, k, v = rearrange(qkv,
23 "n h w (t nh e) -> t n nh (h w) e",
24 t=3, e=self.d_head)
25 q, k = scale(q, k, self.scale[:, None, None], 1e-6)
26 q = apply_rotary_emb(q, theta)
27 k = apply_rotary_emb(k, theta)
28 if self.jk_proj is not None: # Enter into InvFusion block
29 y = self.jnorm(y, cond)
30 jk = self.jk_proj(torch.cat([x, y], -1))
31 jv = self.jv_proj(y)
32 jk, jv = (rearrange(a,
33 "n h w (nh e) -> n nh (h w) e",
34 e=self.d_head) for a in (jk, jv))
35 _, jk = scale(q, jk, self.scale[:, None, None], 1e-6)
36 jk = apply_rotary_emb_(jk, theta)
37 k = torch.cat([k, jk], -2)
38 v = torch.cat([v, jv], -2)
39 x = F.scaled_dot_product_attention(q, k, v, scale=1.0)
40 x = rearrange(x,
41 "n nh (h w) e -> n h w (nh e)",
42 h=skip.shape[-3], w=skip.shape[-2])
43 x = self.dropout(x)
44 x = self.out_proj(x)
45 return x + skip
A.3 Degradations
A.3.1 Patch In-Painting
Patch in-painting is implemented as any in-painting mask where $p\in(0,0.1)$ of patches of some unvarying size are visible. When creating a new mask, the patch size, $p$, and visible patches are randomly sampled until a valid mask in which some patches are visible is created. We perform this “rejection sampling” to remove masks in which no patches are visible.
A.3.2 Strided Motion-Blur
This family of degradations implements a strided convolution operator, for which we specifically choose a motion-blur kernel. To create a new degradation operator, we first sample a motion-blur kernel the size of the input image using the motionblur library. [^4] https://github.com/LeviBorodenko/motionblur We then sample a stride size from $[2,4,8]$ and smooth the motion-blur kernel by convolving it with a isometric Gaussian kernel with standard deviation equal to the stride. The kernels are normalized for numeric stability.
A.3.3 Matrix Degradation
The matrix degradation is a matrix operator with 21282-128 rows, each of which is sampled from a multivariate normal distribution and normalized. The number of rows is limited to enable the use of different matrices across the batch dimension, which is more computationally intensive. The pseudo-inverse operator is computed directly with pytorch’s p_inv function.
A.3.4 Box Out-Painting
A.4 Zero-Shot Methods
@@ Appendix B Using Diffusion Models for MMSE
In Sec. 3.1 we explore the use of the InvFusion architecture of MMSE estimation. A network can be trained directly as a degradation-aware MMSE estimator using a regression loss as seen in Eq.˜7. Nevertheless, we notice that we can utilize our existing trained diffusion models (trained using Eq.˜6) for the same task. Intuitively, this is because for sufficiently high noise values, the noisy input to the network is equivalent to pure noise. Therefore, The network learns to rely only on the information from $\mathbf{y}$} and $\bm{H}$}. In practice, we use the value $\sigma_{t}=100$ for the MMSE estimation.
We test the gap between models trained directly for MMSE estimation (Eq.˜7) and models trained for diffusion (Eq.˜6) on FFHQ64, using the same architecture. The MMSE model’s input is Gaussian noise, instead of the noisy image input to the diffusion model, to keep the architecture identical. The results in Tab.˜7 reflect that both methods yield approximately the same results, whether trained using the degradation-aware InvFusion architecture and loss or without it. We conclude that it is probably more cost-effective to use a single model for both posterior sampling and MMSE estimation task, and do so for the experiments in our paper.
@@
Trained for MMSE (Eq.˜7)
Palette [saharia2022palette] $23.08$ $19.36$ $15.63$
InvFusion $24.66$ $19.79$ $17.92$
Trained For diffusion (Eq.˜6)
Palette [saharia2022palette] $23.47$ $19.65$ $15.51$
InvFusion $24.83$ $19.79$ $17.60$
Appendix C Ablations
Table˜8 shows the effect of applying the InvFusion block only on a fraction of network layers. In this experiment, the InvFusion block has been added or removed selectively to each of the HDiT layers in a specified resolution as used in our base architecture. The results show a gradual improvement across metrics with added applications of the InvFusion block.
Table 8: Ablation of removing the InvFusion block from the architecture at different resolutions.
Strided Motion Blur In-Painting Matrix Operator
[…]
None (Baseline) $21.0$ $4.7$ $5.8$ $17.0$ $5.3$ $5.5$ $12.5$ $6.6$ $25.5$
Resolution 1 $22.3$ $4.5$ $4.5$ $17.1$ $4.9$ $4.9$ $14.6$ $8.9$ $8.9$
Resolutions 1 & 2 $22.7$ $4.4$ $4.4$ $17.1$ $5.0$ $5.0$ $15.0$ $7.9$ $7.9$
Resolutions 1 - 3 (All) $22.7$ $4.2$ $4.2$ $17.1$ $5.0$ $5.0$ $15.1$ $7.1$ $7.1$
Appendix D Additional Results
D.1 LPIPS Evaluation
We include LPIPS [lpips] evaluations for our experiments in Tab.˜9, which correspond to the results shown in Tabs. 1 and 2 in Sec.˜4. InvFusion demonstrated superior performance on LPIPS as well.
Table 9: LPIPS evaluation
Strided Motion Blur In-Painting Matrix Operator
Dataset FFHQ ImageNet FFHQ ImageNet FFHQ ImageNet
Zero-shot Methods
DDRM [kawar2022denoising] 0.095 0.246 0.268 0.458 0.517 0.277
DDNM [wang2022zero] 0.105 0.373 0.171 0.406 0.672 0.300
DPS [chung2023diffusion] 0.136 0.233 0.276 0.462 0.480 0.281
PiGDM [song2023pseudoinverseguided] 0.064 0.156 0.158 0.312 0.470 0.287
DAPS [zhang2024improving] 0.683 0.736 0.148 0.327 0.535 0.263
MGPS [moufad2024variational] 0.250 0.441 0.193 0.371 0.085 0.193
Training-based Methods
Palette [saharia2022palette] 0.068 0.147 0.103 0.215 0.409 0.236
InvFusion 0.053 0.122 0.101 0.213 0.340 0.182
MMSE predictor (Training-based)
Palette [saharia2022palette] 0.126 0.327 0.144 0.318 0.606 0.314
InvFusion 0.106 0.278 0.143 0.315 0.469 0.237
D.2 Consistency Comparison
To validate the consistency of different methods directly, we measure the MSE in the degradation range-space for each reconstruction method. The results shown in Tab.˜10 demonstrate the consistency of the reconstructions corresponding to the experiment in Tab.˜2. MSE is used to accommodate degradations of differing rank and range, and therefore has a different scale for each degradation class. The MSE values confirm that InvFusion achieves consistency comparable to methods that employ explicit projections, such as DDNM or DDRM. These results offer a complementary metric to the FID/CFID analysis.
Table 10: Consistency comparison using MSE in the degradation space.
Method Strided Motion Blur In-Painting Matrix Operator
DDRM [kawar2022denoising] $6.90$ $0.00$ $0.00$
DDNM [wang2022zero] $2.28$ $0.00$ $0.00$
DPS [chung2023diffusion] $789.63$ $10.27$ $0.00$
[…]
DAPS [zhang2024improving] $720765.61$ $2.89$ $5.68$
MGPS [moufad2024variational] $2087.28$ $23.55$ $9.30$
Palette [saharia2022palette] $115.67$ $0.23$ $11.57$
InvFusion $5.94$ $0.01$ $0.00$
D.3 Noisy Degradations
Table 11 shows the results of the experiment described in Sec.˜4.2. The results show that InvFusion maintains superiority when noise is added to the degradation. We expect this framework to work well across many noise types, not limited to white Gaussian noise, as the model learns to treat the noise through training, which does not rely on any fundamental properties of the noise. Nevertheless, for high noise levels InvFusion fails, as seen in the case for high levels of Poisson noise shown in  Tab.˜12.
@@ D.3 Noisy Degradations
DDRM [kawar2022denoising] $11.7$ $137.0$ $12.2$ $155.6$ $11.8$ $87.0$ $25$
DDNM [wang2022zero] $19.8$ $97.2$ $17.3$ $46.4$ $11.9$ $136.9$ $100$
DPS [chung2023diffusion] $7.0$ $12.3$ $8.3$ $7.8$ $9.2$ $22.9$ 10001000*
[…]
DAPS [zhang2024improving] $16.9$ $46.4$ $15.5$ $43.7$ $10.6$ $76.8$ $1000$
@@
Palette [saharia2022palette] $17.1$ $6.0$ $16.7$ $5.2$ $12.6$ $6.6$ $63$
InvFusion $21.7$ 4.3 $16.6$ 5.0 $14.7$ 6.2 $63$
MMSE predictor (Training-based)
Palette [saharia2022palette] $19.8$ - $18.9$ - $15.6$ - 11
InvFusion 24.1 - 19.4 - 17.5 - 11
    Methods that use the network derivative, using more computation and memory per NFE.
Table 12: Comparison of inverse-problem with high level of Poisson noise on FFHQ64.
Strided Motion Blur In-Painting
[…]
Palette [saharia2022palette] $14.1$ $17.6$ $12.6$ $9.9$
InvFusion $11.9$ $60.0$ $14.7$ $11.5$
D.4 FFHQ 256
Table 13 shows quantitative results for our experiment on $256\times 256$ FFHQ images. InvFusion remains the SOTA method as measured by CFID – the perceptual image quality of valid solutions to the inverse problem.
@@ D.4 FFHQ 256
DDRM [kawar2022denoising] $25.3$ $43.0$ $43.0$ $12.8$ $101.6$ $101.6$ $25$
DDNM [wang2022zero] $26.2$ $49.9$ $49.9$ $15.2$ $102.1$ $102.1$ $100$
DPS [chung2023diffusion] $22.6$ $15.2$ $19.7$ $9.8$ $24.3$ $90.7$ 10001000*
[…]
DAPS [zhang2024improving] $9.4$ $397.8$ $145.7$ $17.8$ $40.1$ $46.8$ $1000$
MPGS [moufad2024variational] $26.0$ $26.5$ $25.9$ $18.2$ $14.8$ $15.2$ $1764$
@@
Palette [saharia2022palette] $23.2$ $8.9$ $12.1$ $18.2$ 11.1 $12.4$ $63$
InDI [delbracio2023inversion] $24.1$ $11.3$ $14.2$ $20.6$ $14.0$ $14.1$ $50$
InvFusion $24.4$ 8.4 8.4 $18.5$ $11.2$ 11.2 $63$
MMSE predictor (Training-based)
Palette [saharia2022palette] $25.4$ - - $20.7$ - - 11
InvFusion 26.4 - - 21.1 - - 11
    Methods that use the network derivative, using more computation and memory per NFE.
D.5 Comparison to Algorithm Unrolling
In this experiment, we compare our method to algorithm unrolling [gregor2010learning, sun2016deep, zhang2018ista, aggarwal2018modl, monga2021algorithm]. Using the unconditional model trained for the experiments in Sec.˜4, we unroll the DDRM [kawar2022denoising] algorithm through 10 steps and train this in an end-to-end fashion with MSE loss. Table˜14 shows the results obtained by DDRM before and after the finetuning, compared to InvFusion. The results show that InvFusion is far superior in terms of image quality, and can reach higher fidelity using a single NFE using the MSE estimator method.
Table 14: Comparison of our method with algorithm unrolling.
Strided Motion Blur In-Painting Matrix Operator NFEs
[…]
DDRM (Sampling) $22.2$ $29.5$ $29.5$ $12.2$ $92.0$ $92.0$ $11.0$ $91.4$ $91.4$ $10$
Unrolled DDRM $24.1$ $41.2$ $40.0$ $19.6$ $65.6$ $65.6$ $14.1$ $182.9$ $182.9$ $10$
InvFusion (Sampling) $22.7$ $4.2$ $4.2$ $17.1$ $5.0$ $5.0$ $15.1$ $7.1$ $7.1$ $63$
InvFusion-MSE $25.0$ - - $19.8$ - - $17.9$ - - 11
D.6 Comparison to Deep Image Prior
In this section, we compare our method to a recent Deep Image Prior (DIP) method, ASeq-DIP [alkhouri2024image]. Methods like Deep Image Prior operate at the single-image level, making restoration a computationally prohibitive task for more than a few images. Additionally, these methods typically optimize for different objectives—such as MAP estimation rather than posterior sampling quality—and employ fundamentally different architectures, limiting the conclusiveness of direct comparisons. To provide meaningful context nonetheless, we evaluated PSNR, KID [kid] and LPIPS [lpips] metrics for 1024 images restored using ASeq-DIP [alkhouri2024image], and compare it to our methods in Tab.˜15. The results demonstrate a clear advantage for our model, despite the substantially cheaper evaluation. We use the default hyperparameters as described in the official implementation, and we find that using a larger model or a different number of inner and outer steps has a negligible effect on the results.
Table 15: Comparison of our method with Deep Image Prior.
Strided Motion Blur In-Painting Matrix Operator
[…]
ASeq-DIP [alkhouri2024image] $22.4$ $51.2$ $0.48$ $13.1$ $75.5$ $0.66$ $12.7$ $247.0$ $0.93$
InvFusion $21.1$ $6.3$ $0.13$ $15.5$ $7.6$ $0.24$ $15.2$ $10.0$ $0.34$
InvFusion-MSE $23.7$ $25.71$ $0.33$ $18.44$ $37.9$ $0.39$ $18.1$ $43.8$ $0.50$
D.7 Unconditional Samples
Figures 6 and 6 show qualitative examples for unconditional samples from our model i.e. samples that are not conditioned on a measurement vector $\mathbf{y}$}. The models trained on ImageNet are conditioned on the class label, and use CFG [ho2022classifier] coefficient of 2.0 to enhance the image quality.
Figure 5: Examples of unconditional samples from FFHQ generated using our model.
Figure 6: Examples of class-conditional samples from ImageNet generated using our model. Samples are not conditioned on any measurements $\mathbf{y}$}, and are sampled with CFG of 2.0 on the class conditioning.